Which of the following is true regarding the series ∑nr=1e1−r
Hint: write out the first few terms of the series to understand it better

A. This series is not a geometric series.
B. ∑r=1ne1−r=en−1e−1
C. limn→∞∑r=1ne1−r=ee−1
D. The series is a telescoping series
1
Expert's answer
2020-05-11T10:22:06-0400

A) this series is a geometric series: "a=1;q=e^{-1}" - false

B) For the geometric series "S_n=\frac{a(1-q^n)}{1-q}" therefore "S_n=\frac{1-e^{-n}}{1-e^{-1}}"

C) "\lim\limits_{n\to\infty}S_n=\frac{1}{1-e^{-1}}=\frac{e}{e-1}" - true

D) this series is not a telescoping series